OpenStudy (anonymous):

How do you write an equation in slope-intercept form for(-2,1); m=1

OpenStudy (katrinakaif):

Plug it into the equation: (y-y1) = m (x-x1) as (-2,1) = (x1,y1) and m = 1

OpenStudy (anonymous):

you're just getting people to do your homework, this question is the exact same as the last one you asked. You know these people won't be here when it comes test time

OpenStudy (katrinakaif):

Can you solve that for me Mira? I can guide you through.

OpenStudy (anonymous):

okay so, 1= 1(-2) +b right?

OpenStudy (katrinakaif):

(y- 1) = 1 (x- (-2) y - 1 = 1 (x + 2) y - 1 = 1x + 2 y = 1x +1

OpenStudy (anonymous):

wait what D: I'm so confused.

OpenStudy (katrinakaif):

Haha, okay lets take it step by step.

OpenStudy (katrinakaif):

The equation we plug these in is: (y -y1) = m (x-x1) The given points are: (-2, 1) which substitutes as (x1,y1) If we put it all together, we get: (y - 1) = m (x - (-2) Understood?

OpenStudy (anonymous):

Yes okay..

OpenStudy (anonymous):

equation is y=mx+b << u have (-2,1), which is x=-2, y=1, and m=1.... plug back in--- 1=1(-2)+b, solve for b, u get b=3, then rewrite, y=1x+3

OpenStudy (katrinakaif):

Now lets plug in for m as well, giving us: (y - 1) = 1 (x - (-2)

OpenStudy (katrinakaif):

(y- 1) = 1 (x + 2) is what it becomes when simplified.

OpenStudy (anonymous):

yeah so 1=1x-2?

OpenStudy (anonymous):

im not following ur way, katrina..... u lost me lol

OpenStudy (katrinakaif):

First: You forgot about the "y" before 1 = 1x - 2

OpenStudy (katrinakaif):

Second: Its y - 1 = 1x + 2 (1x PLUS 2)

OpenStudy (anonymous):

(y- 1) = 1 (x- (-2) y - 1 = 1 (x + 2) y - 1 = 1x + 2 <<<sign error y = 1x +1 << should be y=1x+3?

OpenStudy (katrinakaif):

Lol Yep.

OpenStudy (anonymous):

got cha, i think the equation y=mx+b is way easier to remember

OpenStudy (katrinakaif):

Hmm, but you have to begin with : (y-y1) = m (x-x1)

OpenStudy (anonymous):

OH I UNDERSTAND THANK YOU!!!!!!